# magnetohydrodynamics (MHD) — turbulence — diffusion — protoplanetary disks

### v1

@@ Abstract
We perform three-dimensional global non-ideal magnetohydrodynamic simulations of a protoplanetary disk containing the inner dead-zone edge. We take into account realistic diffusion coefficients of the Ohmic resistivity and ambipolar diffusion based on detailed chemical reactions with a single-size dust grains. We found that the conventional dead zone identified by the Elsässer numbers of the Ohmic resistivity and ambipolar diffusion is divided into two regions: ”the transition zone” and ”the coherent zone”. The coherent zone has the same properties as the conventional dead zone, and extends outside of the transition zone in the radial direction. Between the active and coherent zones, we discover the transition zone whose inner edge is identical to that of the conventional dead-zone. The transition zone extends out over the regions where thermal ionization determines diffusion coefficients. The transition zone has completely different physical properties than the conventional dead zone, the so-called undead zone, and zombie zone. Combination of amplification of the radial magnetic field owing to the ambipolar diffusion and a steep radial gradient of the Ohmic diffusivity causes the efficient evacuation of the net vertical magnetic flux from the transition zone within several rotations. Surface gas accretion occurs in the coherent zone but not in the transition zone. The presence of the transition zone prohibits mass and magnetic flux transport from the coherent zone to the active zone. Mass accumulation occurs at both edges of the transition zone as a result of mass supply from the active and coherent zones.
@@ 2.1 Basic Equations
$\frac{\partial\rho\bm{v}}{\partial t}+\bm{\nabla}\cdot\left(\rho\bm{v}\otimes\bm{v}+{\bf T}\right)=-\rho\bm{\nabla}\phi,$ (2)
@@
where $\rho$ is the gas density, $\bm{v}$ is the gas velocity, $\bm{B}$ is the magnetic field, $\phi$ is the gravitational potential of the central star, $\bm{J}$ is the current density, $\bm{J}_{\perp}\equiv\bm{J}-\bm{J}\cdot\bm{B}/|\bm{B}|$ is the electric current density perpendicular to the local magnetic field, $P$ is the gas pressure, ${\bf T}$ is the stress tensor, ${\bf I}$ is the identity matrix, $\eta_{\mathrm{O}}$ and $\eta_{\mathrm{A}}$ are the diffusion coefficients for OR and AD, respectively. They are defined in Section 2.2.2. For simplicity, the Hall effect is neglected in this work.
In this paper, for simplicity we use the locally isothermal equation of state where the gas temperature remains the same as the initial one (Suzuki & Inutsuka, 2014; Zhu & Stone, 2018). To achieve this, we solve the MHD equations including the energy equation with $\gamma=5/3$, and the gas pressure is derived from the locally isothermal condition where the temperature distribution is shown in Equation (10). We note that a locally isothermal equation of state potentially drives the vertical shear instability (VSI) (Nelson et al., 2013). In the active zone, Latter & Papaloizou (2018) shows that the VSI is suppressed by magnetic fields in the ideal MHD limit (also see Cui & Lin, 2021; Latter & Kunz, 2022). In the dead-zone, OR can revive the VSI if the Elässer number is sufficiently low. Latter & Kunz (2022) show that the behaviours of AD are rather complicated because the VSI is coupled with AD shear-instability (Kunz, 2008). In our simulations, at least in the regions around the inner dead-zone edge, which is focused on in this paper, the VSI does not appear to grow because the gas is significantly disturbed by the MRI turbulence in the active zone. The radial velocity dispersion is much larger than the vertical one, which is not expected in the VSI growth mode. By contrast, outer regions far from the inner dead-zone edge can suffer from the VSI. Our simulation time however is too short to see the growth the VSI in such outer regions.
@@ 2.2.1 Initial Condition
[…]
@@ 2.2.2 Diffusion Coefficients
[…]
@@
For the non-thermal ionization, we use the data table of the three conductivities based on a chemical reaction network of ${}^{-}$, ${}^{+}$, ${}^{+}$, ${}^{+}$, H+${}_{3}^{+}$, ${}^{+}$, ${}^{+}$ in the gas phase and the charged dust grains using the methods described by Nakano et al. (2002) and Okuzumi (2009). The table is a function of $\rho/B$, $T$, and $\rho/\zeta$, where $\zeta$ is the ionization rate that will be defined in Equation (20) and $T$ is replaced with $Tf_{\mathrm{boost}}^{-1}$. For simplicity, the single size dust grain with a radius of $a=0.1~{}\mu$m and an internal density of $2~{}\mathrm{g~{}cm^{-3}}$ is considered. In this paper, the dust-to-gas mass ratio is set to $10^{-4}$, assuming that dust growth reduces the amount of dust grains smaller than the interstellar value $\sim 0.01$. The following three kinds of non-thermal ionization sources are taken into account,
@@
[…]
@@
[…]
@@ 2.3.1 Definition of the Dead Zone Boundaries
The non-ideal MHD effects affect the growth rate of the MRI when the dimensionless Elsässer numbers
@@
are smaller than unity, where $v_{\mathrm{A}}=B/\sqrt{4\pi\rho}$ is the Alfvén speed (Sano & Miyama, 1999; Blaes & Balbus, 1994; Wardle et al., 1999; Kunz & Balbus, 2004; Bai & Stone, 2011).
@@ 2.4 Methods
To solve the basic equations (1)-(3), we use Athena++ (Stone et al., 2020) which is a complete rewrite of the Athena MHD code (Stone et al., 2008). The HLLD (Harten-Lax-van Leer-Discontinuities) method is used as the MHD Riemann solver (Miyoshi & Kusano, 2005). Magnetic fields are integrated by the constrained transport method (Evans & Hawley, 1988; Gardiner et al., 2008). The second-order Runge-Kutta-Legendre super-time stepping technique is used to calculate the magnetic diffusion processes (Meyers et al., 2014), where the time step is limited so that it is not greater than 30 times the time step determined by OR and AD.
@@
A spherical-polar coordinate system $(r,\theta,\phi)$ is adopted in the simulation box $r_{\mathrm{in}}\leq r\leq r_{\mathrm{out}}$, $0\leq\theta\leq\pi$, and $0\leq\phi<2\pi$, where $r_{\mathrm{in}}=0.1~{}\mathrm{au}$ and $r_{\mathrm{out}}=12.5~{}\mathrm{au}$. The radial cell width increases with radius by a constant factor so that the radial cell width is almost identical to the zenith one.
@@ 2.6 Normalization Units
[…]
@@ 3.1.1 Active Zone ($R\leq R_{\mathrm{MRI}}$)
The MRI turbulence drives gas accretion in the inner region of the active zone. The outer region expands outward by receiving the angular momentum from the inner region (Lynden-Bell & Pringle, 1974). In the upper atmospheres ($|z|>2H$), the magnetic torque of the coherent magnetic field drives coherent gas motion near the surface layers.
@@ 3.1.2 Transition Zone ($R_{\mathrm{MRI}}<R<R_{\eta}$)
Figure 9a clearly shows that a density peak appears at each edge of the transition zone, or $R\sim R_{\mathrm{MRI}}$ and $R\sim R_{\eta}$. This structure is formed by a combination of no net radial gas motion in the transition zone and the gas supply from the active zone ($R\leq R_{\mathrm{MRI}}$) and the coherent zone ($R\geq R_{\eta}$). Details will be investigated in Sections 4.1 and 4.2.2.
@@ 3.2.1 Ring Formation in the Active Zone
The resolution dependence on the density distributions are shown in the bottom panels of Figure 11. Although the density is lower for HighRes run than for LowRes run, ring structures are also found in HighRes run. This suggests that if the long term simulation performed for HighRes run, the ring structures would develop.
@@
Figure 12: (a) Normalized vertical mass loss rate in the active zone $C_{\mathrm{w}}$ defined in Equation (37) as a function of radius. The solid and dashed lines correspond to the radial profiles of $C_{\mathrm{w}}$ where $C_{\mathrm{w}}>0$ and $C_{\mathrm{w}}<0$, respectively. The thin dashed lines show the radial profiles of $\langle\beta_{z,\mathrm{mid}}\rangle_{t}^{-1}$ whose value is shown on the right vertical axis. The data are averaged over $900\leq t/t_{\mathrm{K0}}\leq 1000$ and $1900\leq t/t_{\mathrm{K0}}\leq 2000$. (b) Color maps of $\langle\rho\rangle$ at $t=2000~{}t_{\mathrm{K0}}$. The black lines show the streamlines of the poloidal magnetic field averaged over $\phi$. The two dotted magenta lines correspond to $z=\pm z_{\mathrm{atm}}$.
@@
Figure 12a shows that in the gap regions where the radial profiles of $\langle\beta_{z,\mathrm{mid}}\rangle_{t}^{-1}$ have local maxima, $C_{\mathrm{w}}$ is negative, indicating that the gas flows into the density gap regions rather than being ejected from the disk in the vertical direction. Why are there no outflows from the gap regions where the magnetic field is strong? Riols & Lesur (2019) pointed out the gas in the gap regions is ejected by ”wind plumes” where the magnetic fields are strong enough to be coherent and are tilted with respect to the $z$-axis (see also Riols & Lesur, 2018). Figure 12b shows that the poloidal magnetic field structures originating from the gap regions do not maintain a large tilt with respect to the $z$-axis as $|z|$ increase, suggesting that the gas is not continuously accelerated along the field lines.
@@
3.2.2 Disk Wind Launched Around the Transition-zone Inner Edge and Non-existence of Gas Accretion in the Transition Zone
The transition zone is disturbed by the influence from the turbulence in the active zone, and exhibits time variations as will be discussed in Sections 3.2.4 and 3.2.5. In this section, we investigate the quasi-steady structure of the transition zone by taking time average.
@@
Around the mid-plane in $0.45~{}\mathrm{au}\lesssim R\lesssim 0.55~{}\mathrm{au}$, hourglass-shaped poloidal magnetic fields develop. The vertical distribution of each component of the magnetic field at $R=R_{\mathrm{MRI}}$ is shown in Figure 14a. The toroidal magnetic field dominates over the other components, and has a sharp gradient at the mid-plane.
@@
AD plays a critical role in the formation of the hourglass-shaped poloidal magnetic fields in $0.45~{}\mathrm{au}\lesssim R\lesssim 0.55~{}\mathrm{au}$. To investigate which of the ideal, OR, and AD terms is the most effective to produce the hourglass-shaped magnetic fields, we measure their contributions to (×$-\langle\left(\nabla\times\bm{E}\right)_{R}\rangle$, which is equal to $\partial\langle B_{R}\rangle/\partial t$, at a given $z$. First, we focus on the region of $|z|\lesssim 0.01~{}\mathrm{au}$ where the mid-plane gas accretion is seen in Figure 13c. Figure 14b shows that the AD term tilts the poloidal magnetic field toward the radial direction of the disk ($\langle(\partial B_{R}/\partial t)_{\mathrm{A}}\rangle>0$ for $z>0$, and $\langle(\partial B_{R}/\partial t)_{\mathrm{A}}\rangle<0$ for $z<0$), and the ideal MHD term does the same. The vertical profile of $\langle B_{R}\rangle$ is almost stationary by the balance between diffusion due to the OR term and amplification due to the AD and ideal MHD terms. Far from the mid-plane ($|z|\gtrsim 0.02~{}\mathrm{au})$, the AD term behaves quite the opposite. The AD term works as diffusion of $\langle B_{R}\rangle$ amplified by the ideal MHD term.
AD also amplifies the toroidal fields near the mid-plane in $0.45~{}\mathrm{au}\lesssim R\lesssim 0.55~{}\mathrm{au}$. Figure 14c is the same as Figure 14b but for $\langle B_{\phi}\rangle$. Figure 14c shows that the signs of $\langle B_{\phi}\rangle$ and $\langle(\partial B_{\phi}/\partial t)_{\mathrm{A}}\rangle$ are the same near the mid-plane ($|z|\lesssim 0.01~{}\mathrm{au}$), indicating that the AD term amplifies $\langle B_{\phi}\rangle$ and steepens its gradient. The vertical profile of $\langle B_{\phi}\rangle$ is kept almost stationary by the OR term smoothing $\langle B_{\phi}\rangle$. The ideal MHD term partially contributes to the steepening of $\langle B_{\phi}\rangle$. In a similar way as $\langle(\partial B_{R}/\partial t)_{\mathrm{A}}\rangle$, far from the mid-plane, the AD term works as diffusion of $\langle B_{\phi}\rangle$ amplified by the ideal MHD term.
The magnetic torque due to $-B_{z}\partial B_{\phi}/\partial z$ drives mid-plane gas accretion (Figures 13a and 13c). Just above the thin mid-plane gas accretion layer, the wind-like gas flows directing outward are driven. This is also caused by the magnetic torque of the coherent magnetic fields (Blandford & Payne, 1982; Bai et al., 2016). The gas streaming lines at lower latitudes return to the mid-plane. As a result, meridional flows are formed in $0.45~{}\mathrm{au}\lesssim R\lesssim 0.5~{}\mathrm{au}$ and $|z|\lesssim 0.05~{}\mathrm{au}$; the streamlines of the gas flows are circulated each in the north and south sides of the disk as shown in Figure 13c.

### v2

@@ Abstract
We perform three-dimensional global non-ideal magnetohydrodynamic simulations of a protoplanetary disk containing the inner dead-zone edge. We take into account realistic diffusion coefficients of the Ohmic resistivity and ambipolar diffusion based on detailed chemical reactions with single-size dust grains. We found that the conventional dead zone identified by the Elsässer numbers of the Ohmic resistivity and ambipolar diffusion is divided into two regions: ”the transition zone” and ”the coherent zone”. The coherent zone has the same properties as the conventional dead zone, and extends outside of the transition zone in the radial direction. Between the active and coherent zones, we discover the transition zone whose inner edge is identical to that of the conventional dead-zone. The transition zone extends out over the regions where thermal ionization determines diffusion coefficients. The transition zone has completely different physical properties than the conventional dead zone, the so-called undead zone, and zombie zone. The combination of amplification of the radial magnetic field owing to the ambipolar diffusion and a steep radial gradient of the Ohmic diffusivity causes the efficient evacuation of the net vertical magnetic flux from the transition zone within several rotations. Surface gas accretion occurs in the coherent zone but not in the transition zone. The presence of the transition zone prohibits mass and magnetic flux transport from the coherent zone to the active zone. Mass accumulation occurs at both edges of the transition zone as a result of mass supply from the active and coherent zones.
@@ 2.1 Basic Equations
$\frac{\partial\rho\bm{v}}{\partial t}+\bm{\nabla}\cdot\left(\rho\bm{v}\otimes\bm{v}+{\bf T}\right)=-\rho\bm{\nabla}\psi,$ (2)
@@
where $\rho$ is the gas density, $\bm{v}$ is the gas velocity, $\bm{B}$ is the magnetic field, $\psi$ is the gravitational potential of the central star, $\bm{J}$ is the current density, $\bm{J}_{\perp}\equiv\bm{J}-\bm{J}\cdot\bm{B}/|\bm{B}|$ is the electric current density perpendicular to the local magnetic field, $P$ is the gas pressure, ${\bf T}$ is the stress tensor, ${\bf I}$ is the identity matrix, $\eta_{\mathrm{O}}$ and $\eta_{\mathrm{A}}$ are the diffusion coefficients for OR and AD, respectively. They are defined in Section 2.2.2. For simplicity, the Hall effect is neglected in this work.
In this paper, for simplicity we use the locally isothermal equation of state where the gas temperature remains the same as the initial one (Suzuki & Inutsuka, 2014; Zhu & Stone, 2018). To achieve this, we solve the MHD equations including the energy equation with $\gamma=5/3$, and the gas pressure is derived from the locally isothermal condition where the temperature distribution is shown in Equation (10). We note that a locally isothermal equation of state potentially drives the vertical shear instability (VSI) (Nelson et al., 2013). In our simulations, the VSI does not grow because of insufficient simulation time as discussed in Section 5.1.3.
@@ 2.2.1 Initial Condition
The values of $\Sigma_{\mathrm{FUV}}$ are highly uncertain, and Perez-Becker & Chiang (2011) found that FUV layers have a thickness of $\sim 0.01-0.1$ g cm-2 in plane-parallel models. We adopt $\Sigma_{\mathrm{FUV}}=0.3~{}\mathrm{g~{}cm^{-2}}$. It is three times larger than the upper limit of that of Perez-Becker & Chiang (2011) because the radial column density is considered instead of the vertical column density (Bai, 2017).
@@ 2.2.2 Diffusion Coefficients
Since the thermal ionization is important in the regions of sufficiently high temperatures, dust grains are assumed to be sublimated in the calculation of chemical reactions. We consider the thermal ionization of potassium, which has an ionization energy of 3.43 eV. The potassium abundance relative to the H2 number density is $3\times 10^{-7}$. The thermal ionization degree $(x_{e})_{\mathrm{T}}$ is derived by the Saha equation where the electron densities are assumed to be equal to the potassium ion densities. As mentioned before, $T$ is replaced with $Tf_{\mathrm{boost}}^{-1}$ in the Saha equation. When the gas temperature exceeds $\sim 1000~{}$K, the thermal ionization of potassium provides a sufficient high ionization and the region becomes MRI active.
@@
For the non-thermal ionization, we use the data table of the three conductivities based on a chemical reaction network of e-, H+, He+, C+, H+${}_{3}^{+}$, HCO+, Mg+ in the gas phase and the charged dust grains using the methods described by Nakano et al. (2002) and Okuzumi (2009). The table is a function of $\rho/B$, $T$, and $\rho/\zeta$, where $\zeta$ is the ionization rate that will be defined in Equation (20) and $T$ is replaced with $Tf_{\mathrm{boost}}^{-1}$. For simplicity, the single size dust grain with a radius of $a=0.1~{}\mu$m and an internal density of $2~{}\mathrm{g~{}cm^{-3}}$ is considered. In this paper, the dust-to-gas mass ratio is set to $10^{-4}$, assuming that dust growth reduces the amount of dust grains smaller than the interstellar value $\sim 0.01$. The following three kinds of non-thermal ionization sources are taken into account,
@@
where $\Sigma_{\mathrm{CR}}=96~{}$g cm-2 (Umebayashi & Nakano, 1981), and $\Sigma_{\theta}^{+}(r,\theta)$ ($\Sigma_{\theta}^{-}(r,\theta)$) is the column density integrated along the $\theta$-direction from $\theta=0$ ($\theta=\pi$) to $\theta$ at a given $r$. Secondly, we consider ionization due to the X-ray radiation from the central star using the fitting formula (Bai & Goodman, 2009, based on the calculation done by Igea & Glassgold (1999)) as follows:
@@
where $\zeta_{1}=6.0\times 10^{-11}$ s-1, $\Sigma_{\mathrm{X,a}}=3.6\times 10^{-3}$ g cm-2, $\alpha=0.4$, and $\Sigma_{r}(\theta,r)$ is the column density integrated along the $r$-direction from the innermost radius of the simulation box to $r$ at a given $\theta$. The third ionization source is provided by radioactive nuclei,
@@ 2.3.1 Definition of the Dead Zone Boundaries
The non-ideal MHD effects weaken the growth rate of the MRI when the dimensionless Elsässer numbers
@@
are smaller than unity, where $v_{\mathrm{A}}=B/\sqrt{4\pi\rho}$ is the Alfvén speed (Sano & Miyama, 1999; Blaes & Balbus, 1994; Wardle, 1999; Kunz & Balbus, 2004; Bai & Stone, 2011).
@@ 2.4 Methods
To solve the basic equations (1)-(3), we use Athena++ (Stone et al., 2020) which is a complete rewrite of the Athena MHD code (Stone et al., 2008). The HLLD (Harten-Lax-van Leer-Discontinuities) method is used as the MHD Riemann solver (Miyoshi & Kusano, 2005). Magnetic fields are integrated by the constrained transport method (Evans & Hawley, 1988; Gardiner & Stone, 2008). The second-order Runge-Kutta-Legendre super-time stepping technique is used to calculate the magnetic diffusion processes (Meyer et al., 2014), where the time step is limited so that it is not greater than 30 times the time step determined by OR and AD.
@@
A spherical-polar coordinate system $(r,\theta,\phi)$ is adopted in the simulation box $r_{\mathrm{in}}\leq r\leq r_{\mathrm{out}}$, $0\leq\theta\leq\pi$, and $0\leq\phi<2\pi$, where $r_{\mathrm{in}}=0.1~{}\mathrm{au}$ and $r_{\mathrm{out}}=12.5~{}\mathrm{au}$. The radial cell width increases with radius by a constant factor so that the radial cell width is almost identical to the meridional one.
@@ 2.6 Normalization Units
In this paper, all the physical quantities are normalized by the following three quantities, the length scale $R_{0}=1~{}\mathrm{au}$, the time scale $t_{0}=\sqrt{R_{0}^{3}/GM_{*}}=0.1~{}$yr, and the surface density $\Sigma_{0}=500$ g cm-2. The gas temperature is normalized by $T_{0}=(\mu m_{\mathrm{H}}/k_{\mathrm{B}})(R_{0}/t_{0})^{2}=6.3\times 10^{5}~{}\mathrm{K}$. Hereafter, we use normalized physical quantities unless otherwise noted except for $R$ and $z$ that are shown in unit of the astronomical unit.
@@ 3.1.1 Active Zone ($R\leq R_{\mathrm{MRI}}$)
The MRI turbulence drives gas accretion in the inner region of the active zone. The outer region expands outward by receiving the angular momentum from the inner region (Lynden-Bell & Pringle, 1974). In the upper atmospheres ($|z|>2H$), the magnetic torque of the coherent magnetic field drives coherent gas motion near the surface layers (Figure 37d).
@@ 3.1.2 Transition Zone ($R_{\mathrm{MRI}}<R<R_{\eta}$)
Figure 9a clearly shows that a density peak appears at each edge of the transition zone, or $R\sim R_{\mathrm{MRI}}$ and $R\sim R_{\eta}$. This structure is formed by a combination of no net radial gas motion in the transition zone and the gas supply from the active zone ($R\leq R_{\mathrm{MRI}}$) and the coherent zone ($R\geq R_{\eta}$). Details will be investigated in Sections 4.1 and 4.2.2. No $B_{z}$ concentration occurs while the density peak forms around the outer-edge of the transition zone (Figure 9d). This is because the magnetic flux drifts outward from the transition zone to the coherent zone as shown in Sections 4.3.2 and 4.3.3.
@@ 3.2.1 Ring Formation in the Active Zone
The resolution dependence on the density distributions are shown in the bottom panels of Figure 11. The density is lower for HighRes run than for LowRes run in the active zone because HighRes run exhibits faster viscous evolution (Figure 36). Despite the difference in the density structures, ring structures are also found in HighRes run. This suggests that if the long term simulation performed for HighRes run, the ring structures would develop.
@@
Figure 12: (a) Normalized vertical mass loss rate in the active zone $C_{\mathrm{w}}$ defined in Equation (37) as a function of radius. The solid and dashed lines correspond to the radial profiles of $C_{\mathrm{w}}$ where $C_{\mathrm{w}}>0$ and $C_{\mathrm{w}}<0$, respectively. The thin dashed lines show the radial profiles of $\langle\beta_{z,\mathrm{mid}}\rangle_{t}^{-1}$ whose value is shown on the right vertical axis. The data are averaged over $900\leq t/t_{\mathrm{K0}}\leq 1000$ and $1900\leq t/t_{\mathrm{K0}}\leq 2000$. (b) Color maps of $\langle\rho\rangle$ averaged over $1900~{}t_{\mathrm{K0}}\leq t\leq 2000~{}t_{\mathrm{K0}}$ The black lines show the streamlines of the poloidal magnetic field averaged over $\phi$. The blue arrows represent the directions of the poloidal velocities averaged over $\phi$. The two dotted magenta lines correspond to $z=\pm z_{\mathrm{atm}}$.
@@
Figure 12a shows that in the gap regions where the radial profiles of $\langle\beta_{z,\mathrm{mid}}\rangle_{t}^{-1}$ have local maxima, $C_{\mathrm{w}}$ is negative, indicating that the gas flows into the density gap regions rather than being ejected from the disk in the vertical direction. The vertical inflow into the density gap regions is evident from the blue arrows in Figure 12b. Why are there no outflows from the gap regions where the magnetic field is strong? Riols & Lesur (2019) pointed out the gas in the gap regions is ejected by ”wind plumes” where the magnetic fields are strong enough to be coherent and are tilted with respect to the $z$-axis (see also Riols & Lesur, 2018). Figure 12b shows that the poloidal magnetic field structures originating from the gap regions do not maintain a large tilt with respect to the $z$-axis as $|z|$ increase, suggesting that the gas is not continuously accelerated along the field lines.
@@
3.2.2 Gas Dynamics Around the Transition-zone Inner Edge and Non-existence of Gas Accretion in the Transition Zone
The transition zone is disturbed by the influence from the turbulence in the active zone, and exhibits time variations as will be discussed in Sections 3.2.4 and 3.2.5. In this section, we investigate the quasi-steady structure of the transition zone by taking time average. We will show that AD plays an important role in forming the hourglass poloidal magnetic fields and driving the mid-plane gas accretion around the inner edge of the transition zone (Figure 6). Disk winds are launched from higher latitudes around the inner edge of the transition zone. In addition, we will also see that almost no gas accretion is driven in the transition zone. We will quantitatively discuss the gas flows driven by the torques in Section 4.1 and 4.2.
@@
Around the mid-plane in $0.45~{}\mathrm{au}\lesssim R\lesssim 0.55~{}\mathrm{au}$, hourglass-shaped poloidal magnetic fields develop. This is also clearly seen as an increasing trend of $B_{R}$ with $z$ for a constant $B_{z}$ across the mid-plane in Figure 14a, which presents the vertical distribution of each component of the magnetic field at $R=R_{\mathrm{MRI}}$. The toroidal magnetic field dominates over the other components, and has a sharp gradient at the mid-plane.
@@
Hourglass-shaped Poloidal Magnetic Field
AD plays a critical role in the formation of the hourglass-shaped poloidal magnetic fields in $0.45~{}\mathrm{au}\lesssim R\lesssim 0.55~{}\mathrm{au}$. To investigate which of the ideal, OR, and AD terms is the most effective to produce the hourglass-shaped magnetic fields, we measure their contributions to (×$-\langle\left(\nabla\times\bm{E}\right)_{R}\rangle$, which is equal to $\partial\langle B_{R}\rangle/\partial t$, at a given $z$. First, we focus on the region of $|z|\lesssim 0.01~{}\mathrm{au}$ where the mid-plane gas accretion is seen in Figure 13c. Since $B_{z}$ is almost constant in $|z|\lesssim 0.01~{}\mathrm{au}$, the mechanism that steepens the $B_{R}$ gradient (Figure 14a) creates an hourglass-shaped magnetic field. Figure 14b shows that the AD term tilts the poloidal magnetic field toward the radial direction of the disk ($\langle(\partial B_{R}/\partial t)_{\mathrm{A}}\rangle>0$ for $z>0$, and $\langle(\partial B_{R}/\partial t)_{\mathrm{A}}\rangle<0$ for $z<0$), and the ideal MHD term does the same. The vertical profile of $\langle B_{R}\rangle$ is almost stationary by the balance between diffusion due to the OR term and amplification due to the AD and ideal MHD terms. Far from the mid-plane ($|z|\gtrsim 0.02~{}\mathrm{au})$, the AD term behaves quite the opposite. The AD term works as diffusion of $\langle B_{R}\rangle$ amplified by the ideal MHD term.
Mid-plane Gas Accretion
The mid-plane accretion flow seen in $0.45~{}\mathrm{au}\lesssim R\lesssim 0.55~{}\mathrm{au}$ (Figure 13c) is driven by the magnetic torque, $-B_{z}\partial B_{\phi}/\partial z$, due to the pinched magnetic field by AD near the mid-plane. Figure 14c shows that the signs of $\langle B_{\phi}\rangle$ and $\langle(\partial B_{\phi}/\partial t)_{\mathrm{A}}\rangle$ are the same near the mid-plane ($|z|\lesssim 0.01~{}\mathrm{au}$), indicating that the AD term amplifies $\langle B_{\phi}\rangle$ and steepens its gradient. The vertical profile of $\langle B_{\phi}\rangle$ is kept almost stationary by the OR term smoothing $\langle B_{\phi}\rangle$. The ideal MHD term partially contributes to the steepening of $\langle B_{\phi}\rangle$. In a similar way as $\langle(\partial B_{R}/\partial t)_{\mathrm{A}}\rangle$, Far from the mid-plane, the AD term works as diffusion of $\langle B_{\phi}\rangle$ amplified by the ideal MHD term.
Disk Wind Launching
Just above the thin mid-plane gas accretion layer, the wind-like gas flows directing outward are driven. This is also caused by the magnetic torque of the coherent magnetic fields (Blandford & Payne, 1982; Bai et al., 2016, see Section 4.1.2). The gas streaming lines at lower latitudes return to the mid-plane. As a result, meridional flows are formed in $0.45~{}\mathrm{au}\lesssim R\lesssim 0.5~{}\mathrm{au}$ and $|z|\lesssim 0.05~{}\mathrm{au}$; the streamlines of the gas flows are circulated each in the north and south sides of the disk as shown in Figure 13c.